# FlexiBit: Fully Flexible Precision Bit-parallel Accelerator Architecture for Arbitrary Mixed Precision AI

### v1

@@ Abstract.
Therefore, in this paper, we propose a new accelerator architecture, FlexiBit, which efficiently supports FP and INT arithmetic in arbitrary precisions and formats. Unlike previous bit-serial designs, which also provide flexibility but at the cost of performance due to its bit-wise temporal processing nature, FlexiBit’s architecture enables bit-parallel processing of any precision and format without compute unit underutilization. FlexiBit’s new capability to exploit non-power of two precision and format led to 1.66$\times$ and 1.62$\times$ higher performance per area on GPT-3 in FP6 targeting a cloud-scale accelerator, compared to a Tensor Core-like architecture and a state-of-the-art bit-parallel flexible precision accelerator, BitFusion, respectively. Also, the bit-parallel nature of FlexiBit’s architecture led to 3.9 $\times$ higher performance per area compared to a state-of-the-art bit-serial architecture, which demonstrates the significant benefits of FlexiBit for emerging precision and formats.
@@
copyright: nonejournalyear: 2024doi: XXXXXXX.XXXXXXXconference: The 52nd IEEE/ACM International Symposium on Computer Architecture; June 21–25, 2025; Tokyo, Japanisbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
The FP format flexibility has a significant impact on the accelerator’s performance and energy. For example, running a model with FP6 (E3M2) quantization on a fixed FP16 processing unit (E5M10) reduces logic utilization by 73% when FP6 data are padded to fit into FP16, as illustrated in Figure 1 (c). Such a challenge is common for fixed point (INT) data types. To eliminate this underutilization, some works approach the problem with bit-serial architectures(Hao et al., 2022),(Jo and Lee, 2020), (Li and Gupta, 2022), (Judd et al., 2016), (Cheng et al., 2024). While bit-serial processing units can adapt to different precisions, their performance is often limited due to the nature of bit-wise temporal processing, which makes it less ideal for recent large models such as LLMs.
Some previous works have proposed bit-parallel architectures that support power-of-two precisions for integer arithmetic (Sharma et al., 2018), (Reggiani et al., 2023), (Hsiao and Chen, 2020), (Fu et al., 2021), (Venkataramani et al., 2021) (Noh et al., 2023a), (Yang et al., 2023), (NVIDIA, 2017). However, none of them support (1) floating-point operations and (2) provide full flexibility of precision for both operands (input and weight).
@@
Therefore, in this work, we explore a new hardware architecture that can support arbitrary bit precision and variants of the FP data formats to eliminate (1) the hardware constraints blocking algorithmic innovations in quantization and (2) the need for replacing hardware for new data precision and formats. One key architecture of our accelerator, FlexiBit, is Flexible Bit Reduction Tree (FBRT) illustrated in Figure 6, which extends the MAERI (Kwon et al., 2018)’s augmented reduction tree toward the bit level. FBRT includes a configurable shifting mechanism and additional links in the tree architecture for flexible precision/format bit-level arithmetic in a bit-parallel manner. Using the flexibility, FlexiBit’s multiplier implemented using FBRT eliminates hardware logic underutilization from data and hardware precision mismatch. We also propose a new flexible-bit adder for exponent addition and data layout management modules, which are crucial for efficient arbitrary precision data processing.
@@
Our evaluation shows that FP format flexibility unlocks significant benefits from non-power-of-two quantization and/or flexible formats. Compared to a Tensor Core (NVIDIA, 2017)-like architecture, FlexiBit provides 59% less latency and 66% less energy, on average when running FP6 arithmetic, across recent language models we evaluate including Bert-base (Kenton and Toutanova, 2019), Llama2-7B (Touvron et al., 2023), Llama-2-70b (Touvron et al., 2023), and GPT-3(Brown et al., 2020). Compared to the state-of-the-art flexible precision architecture, Bit-fusion (Sharma et al., 2018) with our extension for floating point arithmetic, FlexiBit provided 31% less latency and 33% energy, on average, for the same workload. While a design configuration of FlexiBit requires only 0.5% and 1% more hardware and area compared to Tensor Core and Bit-fusion(adopted for FP arithmetic operations), respectively, our analysis of the performance per area shows that the benefits from the extra area surpass the cost, with 20% and 10% higher performance per area, on average for cloud-scale accelerator configurations, which demonstrates that the benefits of FlexiBit are significantly larger than the costs.
@@
    We highlight the challenges from recent non-trivial FP quantization methods to the hardware architecture.
@@
    We develop a flexible FP precision/format accelerator architecture, FlexiBit, which can efficiently support arbitrary FP precision/format and completely eliminate hardware-oriented constraints on FP/INT quantization algorithms.
@@ 2.1. Floating Point Representation and Arithmetic
Figure 2. An example of representing a binary floating point number in IEEE 754 (IEE, 2019) FP16.
Floating-point (FP) arithmetic operations enable the representation and manipulation of real numbers with a high degree of precision and range. Most FP operations in computers adhere to the IEEE 754 standard (IEE, 2019), which defines the encoding of floating-point numbers. The standard encoding style is comprehensive to include arbitrary bit precision with different number of exponent and mantissa bits. To provide an insight into the floating point number representation, we provide an floating-point encoding example in Figure 2.
In this paper, we refer to the different distribution of bits to exponent and mantissa as format. For example, in FP8 precision, we can distribute 4 bits to the exponent and 3 bits to the mantissa (less sign bit) less 1 bit for sign, which constructs the e4m3 format. Alternatively, we can distribute 5 and 2 bits for exponent and mantissa, respectively, which constructs the e5m2 format.  Figure 1 (b) illustrates such variants from FP quantization works (Micikevicius et al., 2022; Xia et al., 2024; Liu et al., 2023).
In deep learning workloads, multiplication and addition are dominant operations (Chen et al., 2016). Therefore, we discuss the multiplication and addition operations with standard FP encoding. Also, we discuss emerging microscaling (MX) format  (Rouhani et al., 2023) arithmetic.
Figure 3. A floating point multiplication example.
Multiplication. Multiplication consists of five steps, as illustrated in Figure 3; (1) adding exponents, (2) determining sign, (3) multiplying mantissas, (4) shift/truncate mantissa to normalize the results for output precision, and (5) adjusting exponent based on the step 4 results. The process of multiplying two floating point numbers A and B can be summarized as follows:
@@
Addition. Because the significand in floating point numbers has different scales based on their exponent values, the addition first requires to align the number scale. The following equations describe the process of adding two floating point numbers, A and B, by shifting A and taking B as the reference scale. We omit the output format conversion for simplicity.
@@
$A+B$ $=A^{\prime}+B=(-1)^{s_{B}}\times(1.0+m^{\prime}_{A}+1.0+0.m_{B})\times 2^{e_{B}-\text{bias}}$
Micro-Scaling (MX) format Arithmetic. Micro-Scaling (MX) data format (Rouhani et al., 2023) is a precision-adaptive representation designed to optimize storage and computational efficiency in machine learning and high-performance computing workloads. It leverages variable bit-widths for encoding values, where the precision adapts based on the range and accuracy requirements of the data. MX achieves significant reductions in memory footprint and bandwidth utilization without substantial losses in computational accuracy. MAC operation of two vectors of $A$ and $W$ in MX-Format follows the following equation:
(1) $C=\text{Dot}(A,W)=X(A)\cdot X(B)\sum_{i=1}^{i=K}\left(P_{i}(A)\times P_{i}(W)\right)$
@@ 2.2. Floating Point Quantization for Language Models
Another trend is the use of various formats within the same bit-width, which assigns different numbers of bits to the exponent and mantissa (e.g., e4m3 and e5m2 for FP8 (Micikevicius et al., 2022), BF16 (Burgess et al., 2019) and FP16 (IEE, 2019)). In addition, a recent work explored a completely flexible format within the precision to provide adaptivity to various dynamic range and precision requirements within a model (Liu et al., 2023). Maintaining the hardware efficiency for such variants is yet another challenge to the hardware since most compute units in current hardware are specialized for specific data types (e.g., TPUv4 supports BF16 and INT8 (Jouppi et al., 2023)), following the principle of specialization for accelerator design. We discuss current hardware support for various precision and formats next.
@@ 2.3. Hardware Support for FP Precision and Formats
As a potential solution, bit-serial accelerators such as Stripes (Judd et al., 2016) provide full flexibility by temporarily processing each bit (i.e., serial manner) in each operand. However, the key limitation of such bit-serial accelerators is the lack of bit-level parallelism, which is disadvantageous for large workloads like LLMs. Also, many bit-serial accelerators are near- or in-memory accelerators(e.g., in-cache accelerator (Eckert et al., 2018)), whose scale is limited to support recent large language models despite their effectiveness on small models. Such limitations of existing hardware architectures for large language models today with various FP precision formats motivate us to design a new hardware architecture satisfying the following requirements:
@@ 3. Flexible Precision Arithmetic Units in FlexiBit
To provide desired performance for recent compute-heavy language models with various FP/INT precision and formats, we aim to (1) design a bit-parallel architecture that performs arithmetic operations in high performance and (2) enable full flexibility for the ALU without underutilizing compute logic for arbitrary FP precision and formats. Accordingly, we develop a new floating point multiplication unit based on our new architecture, flexible bit reduction tree (FBRT), which allows parallel multiplication of arbitrary-length mantissa without compute unit underutilization. FBRT extends the the Augmented Reduction Tree (ART) in MAERI (Kwon et al., 2018) for the bit-level flexibility. FBRT architecture enables aggregation of any bit-level multiplication primitives (i.e., cross-product AND results of each operand bit), which is the key enabler for flexible precision. In addition to the FBRT-based flexible multiplication, FlexiBit’s processing element (PE), illustrated in Figure 4, contains modules for flexible-bit exponent addition, truncation and shifting of bits, exponent normalization, and temporal accumulation of partial outputs (temporal accumulation is activated if required by dataflow). We provide an overview of the PE architecture and discuss each component next.
Figure 4. Processing element (PE) architecture in FlexiBit. Red and blue dotted boxes show datapaths for floating point multiplication and accumulation.
@@ 3.1. PE Architecture Overview
The Processing Element in FlexiBit supports both FP and INT multiplication and addition, as illustrated in Figure 4. We categorize hardware modules into two groups by their usage: One group dedicated to FP multiplication (red dotted box in Figure 4) and the other group for FP accumulation (blue dotted box in Figure 4). The PE architecture has design time configurable design parameters as listed in Table 1. One key parameter that significantly affects the performance and cost is $reg\_width$. $reg\_width$ refers to the bit width of activation and weight registers and dictates the amount of parallelism offered by the PE datapath. Note that FlexiBit’s PE can also handle integer data. For integer arithmetic, we bypass FP-dedicated hardware modules, which include flexible bit exponent adder and exponent normalization. We discuss details of each component following the datapath flow from the input operands to the output next.
Table 1. Processing element design parameters
Parameter Description Default val.
$reg\_width$ Weight/act register bit width 24
$R_{M}$ Mantissa register size in bits 12
$R_{E}$ Exponent register size in bits 12
$R_{S}$ Sign register size in bits 12
$L_{\text{prim}}$ Primitive generator bit width 144
$L_{Add}$ The bit width of the FBEA 144
$L_{Acc}$ The bit width of the Accumulator 144
$L_{CST}$ The bit width of the Concat-shift Tree 144
Table 2. Data type parameters
Parameter Description
$BW_{M}(A)$ Activation mantissa bit width
$BW_{M}(W)$ Weight mantissa bit width
$BW_{E}(A)$ Activation exponent bit width
$BW_{E}(W))$ Weight exponent bit width
$P(A)$ Data precision of activation values
$P(W)$ Data precision of weight values
@@ 3.2. Sign / Exponent / Mantissa Separator
The implication of the challenges is that we need a flexible bit routing mechanism to handle arbitrary cases. Our design strategy is enabling flexible bit routability using small crossbars and restricting the width of processed data at one time (i.e., $register\_width$) to constrain the hardware costs for crossbars. For example, if we adopt $reg\_width=24$ and $R_{M}=12$ and $R_{E}=12$, we need $24\times 12$ crossbars to provide full flexibility. As shown in the example, the size of crossbar switches are determined based on $R_{M}(A)$ and $R_{E}(A))$.
@@
LABEL:lst:separator describes the logic in Primitive Generator using a Python-style pseudo code for activation. In LABEL:lst:separator, act_sign_reg, act_mantissa_reg, and act_exponent_reg refer to the corresponding registers in Figure 4 (blue, red, and green registers). The same logic is applied to weight as well. Note that control signals are all the same across the PEs for the same layer. Therefore, we generate the control signals in a centralized controller and broadcast the control signals to all the PEs.

### v2

@@ Abstract.
Therefore, in this paper, we propose a new accelerator architecture, FlexiBit, which efficiently supports FP and INT arithmetic in arbitrary precisions and formats. Unlike previous bit-serial designs, which also provide flexibility but at the cost of performance due to its bit-wise temporal processing nature, FlexiBit’s architecture enables bit-parallel processing of any precision and format without compute unit underutilization. FlexiBit’s new capability to exploit non-power of two precision and format led to 1.66$\times$ and 1.62$\times$ higher performance per area on GPT-3 in FP6 targeting a cloud-scale accelerator, compared to a Tensor Core-like architecture and a state-of-the-art bit-parallel flexible precision accelerator, BitFusion, respectively. Also, the bit-parallel nature of FlexiBit’s architecture led to 3.9 $\times$ higher performance per area compared to a state-of-the-art bit-parallel architecture, which demonstrates the significant benefits of FlexiBit for emerging precision and formats.
@@
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: The 58th IEEE/ACM International Symposium on Microarchitecture; October 18–22, 2025; Seoul, Koreaisbn: 978-X-XXXX-XXXX-X/XX/XX
@@ 1. Introduction
The FP format flexibility has a significant impact on the accelerator’s performance and energy. For example, running a model with FP6 (E3M2) quantization on a fixed FP16 processing unit (E5M10) reduces logic utilization by 73% when FP6 data are padded to fit into FP16, as illustrated in Figure 1 (c). Such a challenge is common for fixed point (INT) data types. To eliminate this underutilization, some works approach the problem with bit-serial architectures(Hao et al., 2022),(Jo and Lee, 2020), (Li and Gupta, 2022), (Judd et al., 2016), (Cheng et al., 2024). While bit-serial processing can adapt to different precisions, their performance is often limited due to the nature of bit-wise temporal processing, which makes it less ideal for recent large models such as LLMs. Some previous works have proposed bit-parallel architectures that support power-of-two precisions for integer arithmetic (Sharma et al., 2018), (Reggiani et al., 2023), (Hsiao and Chen, 2020), (Fu et al., 2021), (Venkataramani et al., 2021) (Noh et al., 2023a), (Yang et al., 2023), (NVIDIA, 2017). However, none of them support (1) FP operations and (2) provide full flexibility of precision for both input and weight.
@@
Therefore, in this work, we explore a new hardware architecture that can support arbitrary bit precision and FP data format variants to eliminate (1) the hardware constraints blocking algorithmic innovations in quantization and (2) the need for replacing hardware for new data precision and formats. One key architecture of our accelerator, FlexiBit, is Flexible Bit Reduction Tree (FBRT) illustrated in Figure 3 (d), which extends the MAERI (Kwon et al., 2018)’s augmented reduction tree toward the bit level. FBRT includes a configurable shifting mechanism and additional links in the tree architecture for flexible precision/format bit-level arithmetic in a bit-parallel manner. Using the flexibility, FlexiBit’s multiplier implemented using FBRT eliminates hardware logic underutilization from data and hardware precision mismatch. We also propose a new flexible-bit adder for exponent addition and data layout management modules, which are crucial for efficient arbitrary precision data processing.
@@
Our evaluation shows that FP format flexibility unlocks significant benefits from non-power-of-two quantization and flexible formats. Compared to a Tensor Core (NVIDIA, 2017)-like architecture, FlexiBit provides 59% less latency and 66% less energy, on average when running FP6 arithmetic, across recent language models we evaluate including Bert-base (Kenton and Toutanova, 2019), Llama2-7B (Touvron et al., 2023), Llama-2-70b (Touvron et al., 2023), and GPT-3(Brown et al., 2020). Compared to the state-of-the-art flexible precision architecture, Bit-fusion (Sharma et al., 2018) (extended for FP support) with our extension for floating point arithmetic, FlexiBit provided 31% less latency and 33% energy, on average, for the same workload. While a design configuration of FlexiBit requires only 0.5% and 1% more hardware and area compared to Tensor Core and Bit-fusion(adopted for FP arithmetic operations), respectively, our analysis of the performance per area shows that the benefits from the extra area surpass the cost, with 20% and 10% higher performance per area, on average for cloud-scale accelerator configurations, which demonstrates that the benefits of FlexiBit are significantly larger than the costs.
@@
    We highlight the challenges from recent non-trivial FP/INT quantization methods to the hardware architecture.
@@
    We develop a fully flexible FP/INT precision/format accelerator architecture, FlexiBit, which can efficiently support arbitrary FP/INT precision/format and completely eliminate hardware-oriented constraints on FP/INT quantization algorithms.
@@ 2.1. Floating Point Representation and Arithmetic
For instance, in FP8 precision, we can distribute 4 bits to the exponent and 3 bits to the mantissa (less sign bit) less 1 bit for sign, which constructs the e4m3 format. Alternatively, we can distribute 5 and 2 bits for exponent and mantissa, respectively, which constructs the e5m2 format.  Figure 1 (b) illustrates such variants from FP quantization works (Micikevicius et al., 2022; Xia et al., 2024; Liu et al., 2023). Here, we discuss the multiplication and addition operations with standard FP encoding as the dominant operations in ML workloads (Chen et al., 2016).
Multiplication. Multiplication consists of (1) adding exponents, (2) determining sign, (3) multiplying mantissas, (4) shift/truncate mantissa to normalize the results for output precision, and (5) adjusting exponent based on the step 4 results, as the following equation represents:
@@
Addition. The following equations describe the process of adding two floating point numbers, A and B, by shifting A and taking B as the reference scale.
@@
$S_{O}$ $=(s_{A}\oplus s_{B})?((max(m^{\prime}_{A},m_{B})==m^{\prime}_{A})?s_{A}:s_{B}):s_{A}$
$A+B$ $=A^{\prime}+B=(-1)^{s_{O}}\times((-1)^{s_{A}}\times 1.m^{\prime}_{A}+(-1)^{s_{B}})\times 1.m_{B}\times 2^{e_{B}-\text{bias}}$
Micro-Scaling (MX) format Arithmetic. Micro-Scaling (MX) data format (Rouhani et al., 2023) is a precision-adaptive representation designed to optimize storage and computational efficiency in machine learning and high-performance computing workloads. It leverages variable bit-widths for encoding values, where the precision adapts based on the range and accuracy requirements of the data. MAC operation of two vectors of $A$ and $W$ in MX-Format follows the following equation:
$C=\text{Dot}(A,W)=X(A)\cdot X(B)\sum_{i=1}^{i=K}\left(P_{i}(A)\times P_{i}(W)\right)$
@@ 2.2. Floating Point Quantization for Language Models
Another trend is the use of various formats within the same bit-width, which assigns different numbers of bits to the exponent and mantissa (e.g., e4m3 and e5m2 for FP8 (Micikevicius et al., 2022), BF16 (Burgess et al., 2019) and FP16 (IEE, 2019)). In addition, a recent work explored a completely flexible format within the precision to provide adaptivity to various dynamic range and precision requirements within a model (Liu et al., 2023). Maintaining the hardware efficiency for such variants is yet another challenge to the hardware since most compute units in current hardware are specialized for specific data types (e.g., TPUv4 supports BF16 and INT8 (Jouppi et al., 2023)), following the principle of specialization for accelerator design. Multiplications and accumulations in non-power-of-two precisions are dominant in aforementioned works. As activations remain in FP16, weights are usually stored in low-precision data types (e.g., FP6 for FP6-LLM (Xia et al., 2024)). Instead of dequantizing the weights to FP16, which is at the cost of hardware support and latency, weights are multiplied with FP16 activations directly to generate FP20 results assuming FP6 is e2m3 format, and accumulators thus perform FP20 accumulations. We also claim that non-standard precisions will be more widely adopted once flexible hardware like FlexiBit emerges, since it allows more fine-grained quantization design space (trade-off between accuracy/perplexity and efficiency) exploration than power-of-two precisions (e.g., to reduce precision from FP8, we need to jump to FP4, ignoring 5,6, and 7 bit options.). Unfortunately, current hardware lacks flexible precision/format supports, as discussed in Section 1. We discuss the details next.
@@ 2.3. Hardware Support for FP Precision and Formats
As a potential solution, bit-serial accelerators such as Stripes (Judd et al., 2016) provide full flexibility by temporally processing each bit (i.e., serial manner) of operands. Many bit-serial accelerators employ near- or in-memory computing due to the memory microarchitecture is friendly to bit level processing (e.g., in-cache accelerator (Eckert et al., 2018)). Although such solutions provide excellent efficiency, they face a major challenge for heavy workloads such as LLMs due to its limitation originating from memory technology (e.g., scale and clock frequency; DRISA (Li et al., 2017)’s adder runs at 1̃ MHz due to long tRC). In addition, the computation in bit-serial accelerators is slowed down at the scale of bit-width due to its bit-serial processing nature, which is a key challenge for handling recent heavy workloads such as LLMs. Such limitations of existing hardware architectures for large language models today with various FP precision and formats motivate us to design a new hardware architecture satisfying the following requirements:
@@ 3. Flexible Precision Arithmetic Units in FlexiBit
To provide desired performance for recent compute-heavy language models with various FP/INT precision and formats, we aim to (1) design a bit-parallel architecture that performs arithmetic operations in high performance and (2) enable full flexibility for the ALU without underutilizing compute logic for arbitrary FP precision and formats. Accordingly, we develop a new floating point multiplication unit based on our new architecture, flexible bit reduction tree (FBRT), which allows parallel multiplication of arbitrary-length mantissa without compute unit underutilization. FBRT extends the the Augmented Reduction Tree (ART) in MAERI (Kwon et al., 2018) for the bit-level flexibility. Our key novelty beyond the ART is (1) bit-level reconfigurability, unlike that in ART’s is data element level, (2) extended node operations (flexible bit shifting, concatenation, and addition, while ART only performs additions), and (3) our carefully designed efficient control modes for complex node operations.
Figure 2. FlexiBit processing element (PE) architecture
FBRT architecture enables aggregation of any bit-level multiplication primitives (i.e., cross-product AND results of each operand bit), which is the key enabler for flexible precision. In addition to the FBRT-based flexible multiplication, FlexiBit’s processing element (PE), illustrated in Figure 2, contains modules for flexible-bit exponent addition, truncation and shifting of bits, exponent normalization, and temporal accumulation of partial outputs (temporal accumulation is activated if required by dataflow). We provide an overview of the PE architecture and discuss each component next.
Table 1. Data annotation and PE design parameters choices
Annotation Description Parameter Description Val.
$BW_{M}(A)$ Activation mantissa bit width $reg\_width$ Weight/act register bit width 24
$BW_{M}(W)$ Weight mantissa bit width $R_{M}$ Mantissa register bit width 12
$BW_{E}(A)$ Activation exponent bit width $R_{E}$ Exponent register bit width 12
$BW_{E}(W)$ Weight exponent bit width $R_{S}$ Sign register bit width 12
$P(A)$ Data precision of activation values $L_{\text{prim}}$ Primitive generator bit width 144
$P(W)$ Data precision of weight values $L_{Add}$ FBEA bit width 144
$L_{Acc}$ Accumulator bit width 144
$L_{CST}$ Concat-shift Tree bit width 144
Figure 3. A walk-through operation example of FP6 activation and FP5 weight from Bit Packing unit to FBRT. For Separator in (b), we only show the weights since its operation is the same for activation.
@@ 3.1. PE Architecture Overview
The Processing Element in FlexiBit supports both FP and INT multiplication and addition, as illustrated in Figure 2. We categorize hardware modules into two groups by their usage: One group dedicated to FP multiplication (red dotted box in Figure 2) and the other group for FP accumulation (blue dotted box in Figure 2). The PE architecture has design time configurable design parameters as listed in Table 1. One key parameter that significantly affects the performance and cost is $reg\_width$. $reg\_width$ refers to the bit width of activation and weight registers and dictates the amount of parallelism offered by the PE datapath. Note that FlexiBit’s PE can also handle integer data. We highlight hardware components used for both integer/FP data and only for FP in  Figure 2. For integer arithmetic, we bypass FP-dedicated hardware modules, which include flexible bit exponent adder and exponent normalization. We discuss details of each component following the datapath flow from the input operands to the output next.
@@ 3.2. Sign / Exponent / Mantissa Separator
The implication of the challenges is that we need a flexible bit routing mechanism to handle arbitrary cases. Our design strategy is enabling flexible bit routability using small crossbars and restricting the width of processed data at one time (i.e., $register\_width$) to constrain the hardware costs for crossbars. For example, if we adopt $reg\_width=24$ and $R_{M}=12$ and $R_{E}=12$, we need $24\times 12$ crossbars to provide full flexibility. Likewise, the size of crossbar switches are determined based on $R_{M}(A)$ and $R_{E}(A))$. To clarify the operation in the separator module, we provide an example in Figure 3 (b).
@@
LABEL:lst:separator describes the Primitive Generator control logic using a Python-style pseudo code for activation. In LABEL:lst:separator, act_sign_reg, act_mantissa_reg, and act_exponent_reg refer to the corresponding registers in Figure 2 (blue, red, and green registers). The same logic is applied to weight as well. Note that control signals are identical across the PEs for the same layer. Therefore, we broadcast the control signals to all the PEs.